# SSL4RL: Revisiting Self-supervised Learning as Intrinsic Reward for Visual-Language Reasoning

### v1

@@ Abstract
Vision-language models (VLMs) have shown remarkable abilities by integrating large language models with visual inputs. However, they often fail to utilize visual evidence adequately, either depending on linguistic priors in vision-centric tasks or resorting to textual shortcuts during reasoning. Although reinforcement learning (RL) can align models with desired behaviors, its application to VLMs has been hindered by the lack of scalable and reliable reward mechanisms. To overcome this challenge, we propose SSL4RL, a novel framework that leverages self-supervised learning (SSL) tasks as a source of verifiable rewards for RL-based fine-tuning. Our approach reformulates SSL objectives—such as predicting image rotation or reconstructing masked patches—into dense, automatic reward signals, eliminating the need for human preference data or unreliable AI evaluators. Experiments show that SSL4RL substantially improves performance on both vision-centric and vision-language reasoning benchmarks. Furthermore, through systematic ablations, we identify key factors—such as task difficulty, model scale, and semantic alignment with the target domain—that influence the effectiveness of SSL4RL tasks, offering new design principles for future work. We also demonstrate the framework’s generality by applying it to graph learning, where it yields significant gains. SSL4RL establishes a versatile and effective paradigm for aligning multimodal models using verifiable, self-supervised objectives.
@@ 1 Introduction
Vision–Language Models (VLMs) have rapidly advanced multimodal understanding by leveraging the expert-level reasoning capabilities of Large Language Models (LLMs). This synergy has enabled broad applications, from visual question answering to interactive dialogue. Yet the reliance on linguistic priors introduces systematic weaknesses. For vision-centric tasks—where answers must be derived solely from image content, such as classification—VLMs often lag behind specialist vision models (Fu et al., 2025; Tong et al., 2024; Fu et al., 2024). Conversely, in vision–language reasoning tasks, VLMs tend to exploit textual knowledge rather than grounding their reasoning in visual evidence, a tendency amplified in long-form generation (Jian et al., 2025). These limitations underscore the need for training methods that reinforce visual grounding and robust reasoning simultaneously.
Reinforcement learning (RL) has emerged as the dominant paradigm for post-training large models, demonstrating that preference-based signals—collected from humans or distilled from AI feedback—can substantially improve helpfulness and alignment (Ouyang et al., 2022; Rafailov et al., 2023; Bai et al., 2022). More recently, verifier-driven RL has shown particular promise: models trained with automatically checkable rewards achieve striking gains in domains such as mathematics and programming (Le et al., 2022; Shao et al., 2024; Yue et al., 2025). However, these successes expose a fundamental bottleneck: outside domains with explicit programmatic verifiers, scalable and reliable rewards are scarce. In such cases, training pipelines often revert to LLM-as-a-judge heuristics, which are biased, noisy, and prone to adversarial manipulation (Raina et al., 2024; Chen et al., 2024b). This raises a key question: how can we obtain abundant, verifiable reinforcement signals for VLMs in domains where external verifiers are absent?
Self-supervised learning (SSL) offers a natural but underexplored answer. SSL has been central to representation learning across modalities: masked language modeling and next-token prediction for text (Devlin et al., 2019; Lewis et al., 2020; Raffel et al., 2020), contrastive learning and masked autoencoders for vision (Gidaris et al., 2018; Noroozi & Favaro, 2016; Doersch et al., 2015; Chen et al., 2020; He et al., 2020; Grill et al., 2020; Caron et al., 2021; He et al., 2022). The principle is simple yet powerful: perturb data, and require the model to reconstruct or discriminate. Crucially, SSL tasks define intrinsically verifiable targets. Given an image and its rotated variant, the ground-truth angle is unambiguous. Unlike preference-based signals, these targets are properties of the data itself, providing dense, reliable supervision at scale.
@@
In this paper, we propose SSL4RL, a general framework that repurposes SSL tasks as verifiable reward functions for RL-based post-training. Instead of treating SSL solely as a pre-training tool (Tong et al., 2024), we reinterpret it through the lens of RL: corrupted inputs define trajectories, correctness defines rewards, and policy optimization drives updates. SSL4RL requires no human labels, external verifiers, or heuristic judges, yet produces dense and scalable reinforcement signals. Importantly, unlike conventional SSL, SSL4RL emphasizes generating natural language reasoning paths to solve vision tasks, thereby bridging perceptual learning and reasoning alignment.
We systematically evaluate SSL4RL on both vision-centric and multimodal reasoning benchmarks. On ImageNet-1K, SSL4RL significantly improves classification accuracy over the base model. For vision–language reasoning tasks, it delivers consistent gains, with average improvements of 9% on MMBench and 8% on SEED-Bench. A key finding is that the effectiveness of SSL tasks in SSL4RL differs from their role in traditional SSL: contrastive objectives, while dominant in pre-training, show limited benefit unless paired with stronger data augmentation, whereas position prediction—often deemed too trivial for SSL—emerges as highly effective in the SSL4RL setting. The ablation study reveals that the efficacy of an SSL4RL task is strongly influenced by model capacity, its semantic alignment with downstream tasks, and the inherent difficulty of the task itself. These findings offer initial insights into what makes a good SSL4RL task.
We demonstrate that the SSL4RL paradigm generalizes beyond vision by extending it to the graph domain. Through three graph-related SSL tasks—attribute masking, neighbor prediction, and link prediction (Hou et al., 2022; Hu et al., 2020)—we achieve marked improvements on node classification and link prediction benchmarks. These findings highlight SSL4RL as a versatile recipe for extracting verifiable rewards with self-supervised tasks.
@@ 2 Related Work
RL training with external verifiers. Reinforcement learning has become a dominant paradigm for aligning large models. RLHF aligns LLMs with human intent through preference data (Ouyang et al., 2022), while Direct Preference Optimization (DPO) reframes preference learning as a contrastive loss without explicit reward models (Rafailov et al., 2023). Constitutional AI replaces human raters with rule-based AI feedback (Bai et al., 2022). More recently, verifier-driven RL has achieved notable success in domains such as code and math, where correctness can be mechanically checked (Le et al., 2022; Shao et al., 2024). However, in domains lacking such verifiers, many systems fall back on LLM-as-a-judge signals. While convenient, these rewards are biased, noisy, and adversarially manipulable (Raina et al., 2024; Chen et al., 2024b), undermining their reliability. This limitation motivates the search for scalable alternatives where correctness is intrinsic to the data.
@@
RL training with self-reward. Several methods attempt to reduce dependence on external labels or verifiers by enabling models to generate their own training signals. Self-Instruct (Wang et al., 2023b) bootstraps fine-tuning data through synthetic instructions, while STaR (Zelikman et al., 2022) and Reflexion (Shinn et al., 2023) improve reasoning via self-generated rationales and feedback. Self-Consistency (Wang et al., 2023a) aggregates multiple sampled rationales to increase reliability. Reinforced Pre-Training (RPT) scales this idea by using RL objectives such as next-token prediction at the pretraining stage (Liu et al., 2025). While these methods reduce the need for human labels, they still optimize toward approximating original task accuracy, often requiring correctness evaluation or model-judging heuristics. In contrast, our approach seeks verifiable, abundant signals outside the original task by reinterpreting SSL tasks as reinforcement rewards.
@@
Self-supervised learning across modalities. Self-supervision has been the foundation of representation learning across domains. In language, pretext tasks include masked language modeling (Devlin et al., 2019; Lewis et al., 2020; Raffel et al., 2020) and next-token prediction. Variants that mask reasoning steps show promise in enhancing mathematical reasoning (Chen et al., 2024a). In vision, early tasks include rotation (Gidaris et al., 2018), jigsaw (Noroozi & Favaro, 2016), and context prediction (Doersch et al., 2015), while modern SSL emphasizes contrastive learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Caron et al., 2021) and generative masking (He et al., 2022). In multimodal learning, CLIP (Radford et al., 2021) and ALBEF (Li et al., 2021) demonstrate the power of contrastive and distillation objectives for aligning vision and language. Graph SSL extends these principles to structural data, with node/edge masking (Hu et al., 2020; Hou et al., 2022) and contrastive augmentations (You et al., 2020). A unifying property of all these objectives is their intrinsically verifiable targets, making them natural candidates for repurposing as RL rewards.
@@ 3 SSL4RL Framework
  • Rotation Prediction (Gidaris et al., 2018): $c(x,y)$ rotates $x$ by an angle $y$. The reward is $r=\mathbb{1}[\hat{y}=y]$, where $\hat{y}$ is the predicted angle.
  • Jigsaw Puzzles (Noroozi & Favaro, 2016): $c(x,y)$ partitions $x$ into a grid and permutes patches by index $y$. The reward is $r=\mathbb{1}[\hat{y}=y]$, where $\hat{y}$ is the predicted permutation.
  • Contrastive Learning (Chen et al., 2020; Radford et al., 2021): $c(x)$ generates augmented views. The reward $r$ is a binary classification, imitating the InfoNCE similarity score that encourages high similarity for positive pairs and low similarity for negatives.
  • Patch Position Prediction (Doersch et al., 2015): $c(x,y)$ extracts a patch from location $y$. The reward is $r=\mathbb{1}[\hat{y}=y]$, where $\hat{y}$ is the predicted location.
@@
Optimization via GRPO. We adopt Grouped Reinforcement Policy Optimization (GRPO) (Shao et al., 2024), an efficient policy-gradient method designed for large-scale LLM training. Given a reference policy $\pi_{0}$ (e.g., the supervised fine-tuned model before RL), GRPO optimizes $\pi_{\theta}$ by maximizing the following regularized objective:
@@
where $R(\tau)$ is the SSL-derived reward, and $\beta$ controls the strength of KL regularization. The KL penalty prevents divergence from the reference distribution and stabilizes training. GRPO performs updates by sampling rollouts, normalizing rewards across groups, and applying clipped policy-gradient updates similar to PPO (Schulman et al., 2017), but in a manner more compute-efficient for large batch LLM training. All experiments in this work apply the same GRPO configuration across tasks, ensuring that differences in outcomes are attributable to the choice of SSL reward.
@@ 4 What Makes a Good SSL4RL Task?
In this section, we examine the core principles for designing effective SSL4RL tasks through comprehensive experiments on vision-language reasoning (Section 4.1) and vision-centric benchmarks (Section 4.3). Our results indicate that standard SSL strategies do not directly transfer to the SSL4RL setting. In Section 4.2, we present ablation studies on task difficulty scaling, model size, and task combinations to elucidate these design principles further.
4.1 Experiments on Vision-Language Reasoning Tasks
@@
SSL Task Settings. For each SSL pretext task, we implement the following configurations. In Rotation, images are rotated counterclockwise by a randomly selected angle from 0°, 90°, 180°, 270°. In Jigsaw, each image is partitioned into a 2×2 grid, and the patches are randomly shuffled. In Contrastive, we apply the standard augmentation pipeline from Chen et al. (2020), including color jittering, grayscale conversion, Gaussian blur, horizontal flipping, and random resized cropping, each with an application probability of 0.2. In Position, the image is divided into four equal quadrants, and the target is to identify which quadrant (upper-left, upper-right, lower-left, lower-right) contains a specified patch.
@@
Training and Evaluation Settings. Our experiments primarily adopt Qwen-2.5-VL-3B/7B-Instruct (Bai et al., 2025) for their moderate sizes and strong reasoning performance. For GRPO training, we set the group size to be 5, the KL loss coefficient to be 0.01, the entropy loss coefficient to be 0, and context length to be 2048. We train the models on 8xA800 GPUs with batch size of 512. For evaluation, we adopt the third-party evaluation tool VLMEvalKit (Duan et al., 2024), with the default sampling configured with a temperature of 0.01, top-p of 0.001, and top-k of 1.
Benchmarks. We assess our approach on two prominent multi-modal benchmarks designed for evaluating vision-language reasoning capabilities: MMBench (Liu et al., 2024) and SEED-Bench (Li et al., 2023). MMBench provides a diverse testbed with over 3,000 multiple-choice questions spanning 20 distinct ability dimensions. All results on MMBench are reported for the DEV-EN split, showing the average performance per category (detailed per-dimension results are in Appendix A). SEED-Bench is a comprehensive benchmark with human annotations for image and video modalities. For evaluation, we select the 9 core dimensions pertaining to image understanding, which comprise 14,232 examples in total.
Results. The evaluation results on MMBench and SEED-Bench are presented in Table 1 and Table 2, respectively. Overall, the proposed SSL4RL paradigm leads to substantial performance gains on both MMBench and SEED-Bench benchmarks. On average, SSL4RL-tuned models outperform the base model by 7.39% on MMBench and 8.94% on SEED-Bench. Notably, SSL4RL achieves a remarkable improvement of 39.00 percentage points (80.54% vs. 41.54%) on the Relation Reasoning task in MMBench. On the Visual Reasoning task of SEED-Bench, the improvement reaches up to 19.63 percentage points (73.41% vs. 53.78%). These consistent improvements validate the effectiveness of SSL4RL in enhancing vision-language reasoning.
@@
Category Model TU VR SU IId IIn IA IL SR IC Average
Jigsaw 62.73 69.50 72.67 73.07 70.10 74.32 60.74 55.71 46.43 68.24
Position 52.38 70.69 73.56 72.75 62.89 72.51 64.62 55.25 64.20 69.77
Maximal Improvement $\uparrow$ 21.06 $\uparrow$ 19.63 $\uparrow$ 13.30 $\uparrow$ 9.83 $\uparrow$ 5.15 $\uparrow$ 9.64 $\uparrow$ 5.83 $\uparrow$ 4.11 $\uparrow$ 3.68 $\uparrow$ 8.94
@@
Analysis. Our experimental analysis reveals that the Rotation and Position pretext tasks consistently yield the strongest performance gains. The success of the Position task is intuitive, as localizing a patch within the global image compels the model to integrate fine-grained local details with the overall scene layout, fostering integrated spatial understanding. In contrast, the remarkable effectiveness of the Rotation task presents a more nuanced insight. Despite its perceptual simplicity for humans, we find the task poses a considerable challenge to VLMs, as evidenced by the base model’s near-chance accuracy. We interpret this as evidence that rotation prediction facilitates “anti-commonsense” learning. The model’s pre-training instills a strong prior for canonical orientations, e.g., people are typically depicted upright. Rotated images violate this prior, forcing the model to reconcile anomalous inputs with its existing knowledge, thereby sharpening its relational reasoning and visual comprehension.
Conversely, the standard Contrastive task leads to performance degradation on some downstream tasks. We hypothesize that its default augmentations are insufficiently challenging, causing the model to overfit to superficial invariances without learning semantically meaningful structures. Subsequent ablation studies confirm this: employing a more aggressive augmentation strategy recovers significant performance. This finding highlights the critical need to align SSL task difficulty with model capacity to elicit generalizable feature learning, as explored in Section 4.2.2.
Observations. Through a qualitative analysis of model responses, we observe two key improvements attributable to SSL4RL training. (1) Sharper Attention: The trained models exhibit more precise attention alignment with text queries. For instance, when queried about “hair” (Figure 3, Left), the base model’s attention is diffuse, whereas our model accurately focuses on the relevant region. This is not merely an effect of sparsity, as evidenced by the “sky” query (Figure 3, Right): our model attends to the entire sky area, while the base model activates only sparse, scattered pixels. (2) Reduced Language Bias: SSL4RL mitigates over-reliance on linguistic priors, fostering greater dependence on visual evidence. For example, when asked about a chandelier’s color (Appendix Figure 8), the base model defaults to a common-sense response (e.g., a typical decorative color), while our model first localizes the object and then answers based on the actual appearance.
@@
Figure 3: Cross-Attention Heatmap Comparison. More instances are shown in Appendix E.

### v2

@@ Abstract
Vision-language models (VLMs) have shown remarkable abilities by integrating large language models with visual inputs. However, they often fail to utilize visual evidence adequately, either depending on linguistic priors in vision-centric tasks or resorting to textual shortcuts during reasoning. Although reinforcement learning (RL) can align models with desired behaviors, its application to VLMs has been hindered by the lack of scalable and reliable reward mechanisms. To overcome this challenge, we propose SSL4RL, a novel framework that leverages self-supervised learning (SSL) tasks as a source of verifiable rewards for RL-based fine-tuning. Our approach reformulates SSL objectives—such as predicting image rotation or reconstructing masked patches—into dense, automatic reward signals, eliminating the need for human preference data or unreliable AI evaluators. Experiments show that SSL4RL substantially improves performance on both vision-centric and vision-language reasoning benchmarks, with encouraging potentials on open-ended image-captioning tasks. Through systematic ablations, we identify key factors—such as data volume, model scale, model choice, task difficulty, and semantic alignment with the target domain—that influence the effectiveness of SSL4RL tasks, offering new design principles for future work. We also demonstrate the framework’s generality by applying it to graph learning, where it yields significant gains. SSL4RL establishes a versatile and effective paradigm for aligning multimodal models using verifiable, self-supervised objectives. Our implementation is open-sourced at https://github.com/PKU-ML/SSL4RL, with models hosted on HuggingFace PKU-ML/SSL4RL for broader accessibility.
@@ 1 Introduction
Vision–Language Models (VLMs) have rapidly advanced multimodal understanding by leveraging the expert-level reasoning capabilities of Large Language Models (LLMs). This synergy has enabled broad applications, from visual question answering to interactive dialogue. Yet the reliance on linguistic priors introduces systematic weaknesses. For vision-centric tasks—where answers must be derived solely from image content, such as classification—VLMs often lag behind specialist vision models [fu2025hidden, tong2024cambrian, fu2024blink]. Conversely, in vision–language reasoning tasks, VLMs tend to exploit textual knowledge rather than grounding their reasoning in visual evidence, a tendency amplified in long-form generation [jian2025lookagainthinkslowly]. These limitations underscore the need for training methods that reinforce visual grounding and robust reasoning simultaneously.
Reinforcement learning (RL) has emerged as the dominant paradigm for post-training large models, demonstrating that preference-based signals—collected from humans or distilled from AI feedback—can substantially improve helpfulness and alignment [ouyang2022instructgpt, rafailov2023dpo, bai2022constitutional]. More recently, verifier-driven RL has shown particular promise: models trained with automatically checkable rewards achieve striking gains in domains such as mathematics and programming [le2022coderl, deepseekmath2024]. However, these successes expose a fundamental bottleneck: outside domains with explicit programmatic verifiers, scalable and reliable rewards are scarce. In such cases, training pipelines often revert to LLM-as-a-judge heuristics, which are biased, noisy, and prone to adversarial manipulation [raina2024llmjudge, chen2024humansorllmsjudge]. This raises a key question: how can we obtain abundant, verifiable reinforcement signals for VLMs in domains where external verifiers are absent?
@@
Self-supervised learning (SSL) offers a natural but underexplored answer. SSL has been central to representation learning across modalities: masked language modeling and next-token prediction for text [devlin2019bert, lewis2020bart, raffel2020t5], contrastive learning and masked autoencoders for vision [gidaris2018rotnet, noroozi2016jigsaw, doersch2015context, chen2020simclr, he2020moco, grill2020byol, caron2021dino, he2022mae]. The principle is simple yet powerful: perturb data, and require the model to reconstruct or discriminate. Crucially, SSL tasks define intrinsically verifiable targets. Given an image and its rotated variant, the ground-truth angle is unambiguous. Unlike preference-based signals, these targets are properties of the data itself, providing dense, reliable supervision at scale.
In this paper, we propose SSL4RL, a general framework that repurposes SSL tasks as verifiable reward functions for RL-based post-training. Instead of treating SSL solely as a pre-training tool [tong2024cambrian], we reinterpret it through the lens of RL: corrupted inputs define trajectories, correctness defines rewards, and policy optimization drives updates. SSL4RL requires no human labels, external verifiers, or heuristic judges, yet produces dense and scalable reinforcement signals. Importantly, unlike conventional SSL, SSL4RL emphasizes generating natural language reasoning paths to solve vision tasks, thereby bridging perceptual learning and reasoning alignment.
We systematically evaluate SSL4RL on both vision-centric and multimodal reasoning benchmarks. On ImageNet-1K, SSL4RL significantly improves classification accuracy over the base model. For vision–language reasoning tasks, it delivers consistent gains, with average improvements of 9% on MMBench and 8% on SEED-Bench. On the open-ended image-captioning platform CapArena [cheng2025caparena], SSL4RL consistently improves over the base model, with the largest performance gain of 8.14 points. A key finding is that the effectiveness of SSL tasks in SSL4RL differs from their role in traditional SSL: contrastive objectives, while dominant in pre-training, show limited benefit unless paired with stronger data augmentation, whereas position prediction—often deemed too trivial for SSL—emerges as highly effective in the SSL4RL setting. The ablation study reveals that the efficacy of an SSL4RL task is influenced by model capacity, its semantic alignment with downstream tasks, and the inherent difficulty of the task itself. These findings offer initial insights into what makes a good SSL4RL task.
We demonstrate that the SSL4RL paradigm generalizes beyond vision by extending it to the graph domain. Through three graph-related SSL tasks—attribute masking, neighbor prediction, and link prediction [hou2022graphmae, hu2019pretrainGNN]—we achieve marked improvements on node classification and link prediction benchmarks. These findings highlight SSL4RL as a versatile recipe for extracting verifiable rewards with self-supervised tasks.
@@ 2 Related Work
RL training with external verifiers. Reinforcement learning has become a dominant paradigm for aligning large models. RLHF aligns LLMs with human intent through preference data [ouyang2022instructgpt], while Direct Preference Optimization (DPO) reframes preference learning as a contrastive loss without explicit reward models [rafailov2023dpo]. Constitutional AI replaces human raters with rule-based AI feedback [bai2022constitutional]. More recently, verifier-driven RL has achieved notable success in domains such as code and math, where correctness can be mechanically checked [le2022coderl, deepseekmath2024]. However, in domains lacking such verifiers, many systems fall back on LLM-as-a-judge signals. While convenient, these rewards are biased, noisy, and adversarially manipulable [raina2024llmjudge, chen2024humansorllmsjudge], undermining their reliability. This limitation motivates the search for scalable alternatives where correctness is intrinsic to the data.
@@
RL training with self-reward. Several methods attempt to reduce dependence on external labels or verifiers by enabling models to generate their own training signals. Self-Instruct [wang2022selfinstruct] bootstraps fine-tuning data through synthetic instructions, while STaR [zelikman2022star] and Reflexion [shinn2023reflexion] improve reasoning via self-generated rationales and feedback. Self-Consistency [wang2023selfconsistency] aggregates multiple sampled rationales to increase reliability. Reinforced Pre-Training (RPT) scales this idea by using RL objectives such as next-token prediction at the pretraining stage [liu2025rpt]. While these methods reduce the need for human labels, they still optimize toward approximating original task accuracy, often requiring correctness evaluation or model-judging heuristics. In contrast, our approach seeks verifiable, abundant signals outside the original task by reinterpreting SSL tasks as reinforcement rewards.
@@
Self-supervised learning across modalities. Self-supervision has been the foundation of representation learning across domains. In language, pretext tasks include masked language modeling [devlin2019bert, lewis2020bart, raffel2020t5] and next-token prediction. Variants that mask reasoning steps show promise in enhancing mathematical reasoning [chen2024maskedthought]. In vision, early tasks include rotation [gidaris2018rotnet], jigsaw [noroozi2016jigsaw], and context prediction [doersch2015context], while modern SSL emphasizes contrastive learning [chen2020simclr, he2020moco, grill2020byol, caron2021dino] and generative masking [he2022mae]. In multimodal learning, CLIP [radford2021clip] and ALBEF [li2021albef] demonstrate the power of contrastive and distillation objectives for aligning vision and language. Graph SSL extends these principles to structural data, with node/edge masking [hu2019pretrainGNN, hou2022graphmae] and contrastive augmentations [you2020graphcl]. A unifying property of all these objectives is their intrinsically verifiable targets, making them natural candidates for repurposing as RL rewards. The most relevant work is Jigsaw-R1 [wang2025jigsawr1], which first establishes jigsaw puzzles as an effective pretext task for MLLM reinforcement learning. We generalize this concept into a unified SSL4RL framework beyond one specific task and can be applied to other domains like graphs. Through comprehensive studies, we systematically investigate the scaling of data volume, model size, task difficulty, and task combinations, offering broader insights into this learning paradigm.
@@ 3 SSL4RL Framework
  • Rotation Prediction [gidaris2018rotnet]: $c(x,y)$ rotates $x$ by an angle $y$. The reward is $r=\mathbb{1}[\hat{y}=y]$, where $\hat{y}$ is the predicted angle.
  • Jigsaw Puzzles [noroozi2016jigsaw]: $c(x,y)$ partitions $x$ into a grid and permutes patches by index $y$. The reward is $r=\mathbb{1}[\hat{y}=y]$, where $\hat{y}$ is the predicted permutation.
  • Contrastive Learning [chen2020simclr, radford2021clip]: $c(x)$ generates augmented views. The reward $r$ is a binary classification, imitating the InfoNCE similarity score that encourages high similarity for positive pairs and low similarity for negatives.
  • Patch Position Prediction [doersch2015context]: $c(x,y)$ extracts a patch from location $y$. The reward is $r=\mathbb{1}[\hat{y}=y]$, where $\hat{y}$ is the predicted location.
@@
Optimization via GRPO. We adopt Grouped Reinforcement Policy Optimization (GRPO) [deepseekmath2024], an efficient policy-gradient method designed for large-scale LLM training. Given a reference policy $\pi_{0}$ (e.g., the supervised fine-tuned model before RL), GRPO optimizes $\pi_{\theta}$ by maximizing the following regularized objective:
@@
where $R(\tau)$ is the SSL-derived reward, and $\beta$ controls the strength of KL regularization. The KL penalty prevents divergence from the reference distribution and stabilizes training. GRPO performs updates by sampling rollouts, normalizing rewards across groups, and applying clipped policy-gradient updates similar to PPO [schulman2017proximal], but in a manner more compute-efficient for large batch LLM training. All experiments in this work apply the same GRPO configuration across tasks, ensuring that differences in outcomes are attributable to the choice of SSL reward.
@@ 4 What Makes a Good SSL4RL Task?
In this section, we examine the core principles for designing effective SSL4RL tasks through comprehensive experiments on vision-language reasoning (Section 4.1), open-ended image-captioning (Section 4.2), and vision-centric benchmarks (Section 4.3). Our results indicate that standard SSL strategies do not directly transfer to the SSL4RL setting. We provide a robustness analysis in Section 4.4. In Section 4.5, we present ablation studies on data volume, model size, base model choice, task difficulty scaling, and task combinations to elucidate these design principles further.
4.1 Vision-Language Reasoning Tasks
@@
SSL Task Settings. For each SSL pretext task, we implement the following configurations. In Rotation, images are rotated counterclockwise by a randomly selected angle from 0°, 90°, 180°, 270°. In Jigsaw, each image is partitioned into a 2×2 grid, and the patches are randomly shuffled. In Contrastive, we apply the standard augmentation pipeline from chen2020simclr, including color jittering, grayscale conversion, Gaussian blur, horizontal flipping, and random resized cropping, each with an application probability of 0.2. In Position, the image is divided into four equal quadrants, and the target is to identify which quadrant (upper-left, upper-right, lower-left, lower-right) contains a specified patch.
@@
Training and Evaluation Settings. Our experiments primarily adopt Qwen-2.5-VL-3B/7B-Instruct [bai2025qwen2] for their moderate sizes and strong reasoning performance. For GRPO training, we set the group size to be 5, the KL loss coefficient to be 0.01, the entropy loss coefficient to be 0, and the context length to be 2048. We train the models on 8xA800 GPUs with a batch size of 512. For evaluation, we adopt the third-party evaluation tool VLMEvalKit [duan2024vlmevalkit], with the default sampling configured with a temperature of 0.01, top-p of 0.001, and top-k of 1. We visualize the rewards, entropy, and response length trajectories during RL training in Appendix K.
Benchmarks. We assess our approach on six prominent Vision Questioning Answer (VQA) benchmarks: MMBench [liu2024mmbench], SEED-Bench [li2023seed], V* [wu2024v], RealWorldQA [xai2024grok15v], BLINK [fu2024blink], and MME-RealWorld-Lite [zhang2024mme], featuring challenging VLMs’ ability on visual perception, spatial understanding, detail capturing, real-world applications, and so on. Detailed datasets descriptions are provided in Appendix I.
Results. Overall, the proposed SSL4RL paradigm leads to substantial performance gains on vision questioning answer tasks. Shown in Table 1 and Appendix Table 23, on average, SSL4RL models outperform the base model by 7.39% on MMBench and 8.94% on SEED-Bench. Notably, SSL4RL achieves a remarkable improvement of 39.00 percentage points (80.54% vs. 41.54%) on the Relation Reasoning task in MMBench. On the Visual Reasoning task of SEED-Bench, the improvement reaches up to 19.63 percentage points (73.41% vs. 53.78%). Results of other benchmarks are presented in Appendix Table 24, where the SSL4RL strategy brings significant improvements, especially on V* (+8.90%) and RealWorldQA (+9.55%). These consistent improvements validate the effectiveness of SSL4RL in enhancing vision-language reasoning. To contextualize the performance of our SSL4RL method, we include a strong baseline VLM-R1 [shen2025vlm]. We fine-tune the base model on 60% of the benchmark and evaluate on the held-out test set. The RL reward is computed directly based on the golden answer. We denote the tuned model as Golden-3B (see more details in Appendix G). As shown in Appendix Table 20 and 21, the performance gap between our SSL4RL variants and the Golden-3B oracle is relatively small (e.g., 81.35% vs. 84.93% on MMBench, and 69.80% vs 73.21% on SEED-Bench), compared to our improvements over base models. This demonstrates that our self-supervised objectives, which require no labeled downstream data, can effectively close the gap to the performance driven by idealized, task-specific reward signals.
@@
Analysis. Our experimental analysis reveals that the Rotation and Position pretext tasks consistently yield the strongest performance gains. The success of the Position task is intuitive, as localizing a patch within the global image compels the model to integrate fine-grained local details with the overall scene layout, fostering integrated spatial understanding. In contrast, the remarkable effectiveness of the Rotation task presents a more nuanced insight. Despite its perceptual simplicity for humans, we find the task poses a considerable challenge to VLMs, as evidenced by the base model’s near-chance accuracy. We interpret this as evidence that rotation prediction facilitates “anti-commonsense” learning. The model’s pre-training instills a strong prior for canonical orientations, e.g., people are typically depicted upright. Rotated images violate this prior, forcing the model to reconcile anomalous inputs with its existing knowledge, thereby sharpening its relational reasoning and visual comprehension. Conversely, the standard Contrastive task leads to performance degradation on some downstream tasks. We hypothesize that its default augmentations are insufficiently challenging, causing the model to overfit to superficial invariances without learning semantically meaningful structures. Subsequent ablation studies confirm this: employing a more aggressive augmentation strategy recovers significant performance. This finding highlights the critical need to align SSL task difficulty with model capacity to elicit generalizable feature learning, as explored in Section 4.5.2. Through a sub-task specific analysis, we find that Rotation exhibits superior performance in Relation Reasoning and Cross-Instance Perception, which can be attributed to its inherent requirement for a structural comprehension of object orientation and spatial relationships. Position demonstrates leading results in Logical Reasoning and holistic Scene Understanding, as its objective of reconstructing a coherent scene from disparate patches fosters robust integrative scene modeling. Jigsaw shows consistent utility in domains reliant on contextual reasoning. Limited by the insufficient augmentations, Contrastive’s optimization for global image similarity appears to inadequately cultivate the fine-grained and relational reasoning capabilities.
@@
Figure 3: Cross-Attention Heatmap Comparison. More instances are shown in Appendix L.
Observations. Through a qualitative analysis of model responses, we observe two key improvements attributable to SSL4RL training. (1) Sharper Attention: The trained models exhibit more precise attention alignment with text queries. For instance, when queried about “hair” (Figure 3), the base model’s attention is diffuse, whereas our model accurately focuses on the relevant region. This is not merely an effect of sparsity, as evidenced by the “sky” query (Appendix Figure 13): our model attends to the entire sky, while the base model activates only scattered pixels. (2) Reduced Language Bias: SSL4RL mitigates over-reliance on linguistic priors, fostering greater dependence on visual evidence. For example, when asked about a chandelier’s color (Appendix Figure 12), the base model defaults to a common-sense response (e.g., a typical decorative color), while our model localizes the object and answers based on the actual appearance.